# PartEdit: Fine-Grained Image Editing using Pre-Trained Diffusion Models

### v1

@@ Abstract
We present the first text-based image editing approach for object parts based on pre-trained diffusion models. Diffusion-based image editing approaches capitalized on the deep understanding of diffusion models of image semantics to perform a variety of edits. However, existing diffusion models lack sufficient understanding of many object parts, hindering fine-grained edits requested by users. To address this, we propose to expand the knowledge of pre-trained diffusion models to allow them to understand various object parts, enabling them to perform fine-grained edits. We achieve this by learning special textual tokens that correspond to different object parts through an efficient token optimization process. These tokens are optimized to produce reliable localization masks at each inference step to localize the editing region. Leveraging these masks, we design feature-blending and adaptive thresholding strategies to execute the edits seamlessly. To evaluate our approach, we establish a benchmark and an evaluation protocol for part editing. Experiments show that our approach outperforms existing editing methods on all metrics and is preferred by users $77-90\%$ of the time in conducted user studies.
@@
Figure 1: Our approach, PartEdit, enables a wide range of fine-grained edits, allowing users to create highly customizable changes. The edits are seamless, precisely localized, and of high visual quality with no leakage into unedited regions.
footnotetext: Project page: https://partedit.github.io/PartEdit/
1 Introduction
@@
Figure 2: A visualization for the cross-attention maps of SDXL [25] that corresponds to different words of the textual prompt. Object parts such as “head” and “hood” are not well-localized, indicating that the model lacks a sufficient understanding of these parts.
@@
Diffusion models [28, 27, 30, 25, 8] have significantly advanced image generation, achieving unprecedented levels of quality and fidelity. This progress is generally attributed to their large-scale training on the LAION-5B dataset [31] with image-text pairs, leading to a profound understanding of images and their semantics. Recent image editing methods [12, 4, 3, 24, 35, 17, 14] have capitalized on this understanding to perform a wide range of edits to enhance the creative capabilities of artists and designers. These methods allow users to specify desired edits through text prompts, enabling both semantic edits, such as modifying objects or their surroundings, and artistic adjustments, like changing style and texture. Ideally, these edits must align with the requested textual prompts while being seamlessly integrated and accurately localized in the image.
@@
In this paper, we address the limitations of pre-trained diffusion models in their understanding of object parts. By expanding their semantic knowledge, we enable fine-grained image edits, providing creators with greater control over their images. We achieve this by training part-specific tokens that specialize in localizing the editing region at each denoising step. Based on this localization, we develop feature blending and adaptive thresholding strategies that ensure seamless and high-quality editing while preserving the unedited areas. Our novel feature blending happens at each layer, at each timestep using nonbinary masks. To learn the part tokens, we design a token optimization process [38] tailored for fine-grained editing, utilizing existing object part datasets [6, 10] or user-provided datasets. This optimization process allows us to keep the pre-trained diffusion model frozen, thereby expanding its semantic understanding without compromising generation quality or existing knowledge. To evaluate our approach and to facilitate the development of future fine-grained editing approaches, we introduce a benchmark and an evaluation protocol for part editing. Experiments show that our approach outperforms all methods in comparison on all metrics and is preferred by users $77-90\%$ of the time in conducted user studies.
2 Text-to-Image Diffusion Models
Diffusion models are probabilistic generative models that attempt to learn an approximation of a data distribution $p(\mathbf{x})$. This is achieved by progressively adding noise to each data sample ${\mathbf{x}_{0}\sim p(\mathbf{x})}$ throughout $T$ timesteps until it converges to an isotropic Gaussian distribution as $T\rightarrow\infty$. During inference, a sampler such as DDIM [32] is used to reverse this process starting from Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ that is iteratively denoised until we obtain a noise-free sample $\hat{\mathbf{x}}_{0}^{0}$. The reverse process for timestep $t\in[T,0]$ is computed as:
$x_{t-1}=$ $\sqrt{\alpha_{t-1}}\ \hat{x}_{0}^{t}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\ \epsilon^{t}_{\theta}(x_{t})+\sigma_{t}\epsilon_{t}\enspace,$ (1)
@@ 2 Text-to-Image Diffusion Models
To generate an image that adheres to a user-provided input, e.g., a textual prompt $\mathcal{P}$, the model can be trained conditionally. In this setting, the UNet is conditioned on the prompt $\mathcal{P}$, and the noise prediction in Equation 1 is computed as $\epsilon^{t}_{\theta}(x_{t},\mathcal{P})$. More specifically, the textual prompt is embedded through a textual encoder, e.g., CLIP [26], to obtain an embedding $E\in\mathrm{R}^{77\times 2048}$ (for SDXL). This textual embedding interacts with the UNet image features $F$ within cross-attention modules at UNet block $i$ to compute attention as:
$A_{i}=\text{{Softmax}}$ $\left(\dfrac{Q_{i}\ K_{i}^{\top}}{\sqrt{{d_{k_{i}}}}}\right)\enspace,$ (2)
@@
3 PartEdit: Fine-Grained Image Editing
Figure 3: An overview of our proposed approach for fine-grained part editing. For an object part $p$, we collect a dataset of images $\mathcal{I}^{p}$ and their corresponding part annotation masks $\mathcal{Y}^{p}$. To optimize a textual token to localize this part, we initialize a random textual embedding $\hat{E}$ that initially generates random cross-attention maps. During optimization, we invert images in $\mathcal{I}^{p}$ and optimize the part token so that the cross-attention maps at different layers and timesteps match the part masks in $\mathcal{Y}^{p}$. After optimizing the token, it can be used during inference to produce a localization mask at each denoising step. These localization masks are used to perform feature bending between the source and the edit image trajectories. Note that we visualize three instances of SDXL [25] for illustration, but in practice, this is done with the same model in a batch of three.
A common approach for editing images in diffusion-based methods involves manipulating the cross-attention maps [12, 24, 7]. If the cross-attention maps do not accurately capture the editing context, e.g., for editing object parts, these methods are likely to fail, as demonstrated in Figure 1. An intuitive solution to this problem is expanding the knowledge of the pre-trained diffusion models to understand object parts. This can be accomplished by fine-tuning the model with additional data of image/text pairs where the text is detailed. However, this approach can be costly due to the extensive annotation required for fine-tuning the model, and there is no guarantee that the model will automatically learn to identify object parts effectively from text.
An alternative approach is leveraging token optimization [38] to learn new concepts through explicit supervision of cross-attention maps. This was proven successful in several applications [11, 18, 22] as it allows learning new concepts while keeping the model’s weights frozen. We leverage token optimization to perform fine-grained image editing of various object parts. We focus on optimizing tokens that can produce reliable non-binary blending masks at each diffusion step to localize the editing region. We supervise the optimization using either existing parts datasets such as PASCAL-Part [6], and PartImageNet [10] or a few user-annotated images.
3.1 Learning Part Tokens
@@
3.2 Choosing Timesteps and UNet Blocks to Optimize
Ideally, we would like to optimize over all timesteps and UNet layers. However, this is computationally and memory-expensive due to the large dimensionality of intermediate features in diffusion models, especially SDXL [25] that we employ. Therefore, we need to select a subset of layers and timesteps to achieve a good balance between localization accuracy and efficiency.
@@ 3.2 Choosing Timesteps and UNet Blocks to Optimize
Figure 4: The impact of the choice of which timesteps to include in the token optimization process. Intermediate timesteps achieve reasonable localization for both big and small parts.
3.3 Part Editing
@@
$\mathcal{T}(X)$ =$=\begin{cases}1&\text{if }X\geq\omega\\ x&\text{if }k/2\leq X<\omega\\ 0,&X<k/2\\ \end{cases},$ (5)
@@
where OTSU is the OTSU thresholding [23], $\omega$ is a tunable tolerance for the transition between the edited parts and the original object. We find that $\omega=3k/2$ achieves the best visual quality, and we fix it for all experiments. This criterion ensures suppressing the background noise and a smooth transition between the edited part and the rest of the object. Finally, we employ $M_{t}$ to blend the features between the source and the editing paths as:
$\hat{F}_{i,t}^{e}=\mathcal{T}(M_{t})\ \hat{F}_{i,t}^{e}+(1-\mathcal{T}(M_{t}))\ \hat{F}_{i,t}^{s}$ (6)
@@
3.4 Real Image Editing
Our approach can also perform part edits on real images by incorporating a real image inversion method, e.g., Ledits++ [3] or EF-DDPM [15]. In this setting, the role of the inversion method is to estimate the diffusion trajectory $x_{0}\dots x_{T}$ for a given real image. This estimated trajectory is then used as the source path in Figure 3. To obtain the source prompt $\mathcal{P}^{s}$ of the real images, we use the image captioning approach, BLIP2 [20], which is commonly used for this purpose. Finally, the edit is applied where the localization and editing paths are similar to the synthetic setting.
3.5 Discussion
@@
Figure 5: A qualitative comparison on synthetic images from the PartEdit benchmark. Our method outperforms both iP2P [4] and P2P [12] on the synthetic setting. Showcasing good localization while integrating seamlessly into the scene, illustrated by the third row with a formal torso edit.
@@
Figure 6: A qualitative comparison against EF-DDPM [15] and Ledits++ [3] on real image editing.
@@
Figure 7: Visualization of editing 2 parts at the same time. PartEdit is able to edit two regions at the same time with minimal modifications. Note that the attention maps showcase average cross-attention across all time steps and layers, while we blend at each timestep, at each layer.
Figure 8: A comparison on synthetic benchmark against Latent Blending [2] and SDXL Inpainting [25] using ground truth masks (✓) against our predicted masks (✗). We observe PartEdit outperforming Latent Blending, which fails totally in some of the edits (spiderman, hair, and chair), while others produce unintegrated edits such as the bear head.
Figure 9: Different edits per image on real image editing using our method. We showcase the versatility of our method, as there is no change in the underlying model; we can leverage the full capabilities of the model without any retraining or fine-tuning. More in Appendix F.
4 Experiments
@@
In this section, we provide an evaluation of our proposed approach, a comparison against existing text-based editing approaches, and an ablation study. To facilitate these aspects, we create a manually annotated benchmark of several object parts. For a comprehensive evaluation, we compare against both synthetic and real image editing approaches.
4.1 Evaluation Setup
@@
Synthetic Image Editing We base our method on a pre-trained SDXL [25]  [^1] Our code, benchmarks, and annotated dataset will be made public. . For sampling, we employ the DDIM sampler [32] with $T=50$ denoising steps and default scheduling parameters as [12]. For token optimization, we train on 10-20 images per token for 2000 optimization steps. We set $t_{e}=T=50$ for complete preservation of the unedited regions. We compare against two popular text-based editing approaches: Prompt-to-Prompt (P2P) [12] and Instruct-Pix2Pix (iP2P) [4]. We use the SDXL implementation of P2P [^2] https://github.com/RoyiRa/prompt-to-prompt-with-sdxl, and the Diffusers implementation of iP2P [^3] https://huggingface.co/docs/diffusers/en/api/pipelines/pix2pix.
@@
Real Image Editing We use Ledits++ [3] inversion method to invert the images where the source prompt $\mathcal{P}^{s}$ is produced by BLIP2 [20] as described in Section 3.4. We compare against Ledits++ and EF-DDPM [15] with the same base model of SD2.1 [28].
@@
PartEdit Benchmark We create a synthetic and real benchmark of 7 object parts: <humanoid-head>, <humanoid-torso>, <human-hair> <animal-head>, <car-body>, <car-hood>, and <chair-seat>. For the synthetic part, we generate random source prompts of the objects of interest at random locations and select random edits from pre-defined lists. We generate a total of 60 synthetic images and manually annotate the part of interest. For the real part, we collect 13 images from the internet and manually annotate and assign editing prompts to them. We denote the synthetic and real benchmarks as PartEdit-Synth and PartEdit-Real, respectively. More details are provided in the Appendix I.

### v2

@@ Abstract.
We present the first text-based image editing approach for object parts based on pre-trained diffusion models. Diffusion-based image editing approaches capitalized on the deep understanding of diffusion models of image semantics to perform a variety of edits. However, existing diffusion models lack sufficient understanding of many object parts, hindering fine-grained edits requested by users. To address this, we propose to expand the knowledge of pre-trained diffusion models to allow them to understand various object parts, enabling them to perform fine-grained edits. We achieve this by learning special textual tokens that correspond to different object parts through an efficient token optimization process. These tokens are optimized to produce reliable localization masks at each inference step to localize the editing region. Leveraging these masks, we design feature-blending and adaptive thresholding strategies to execute the edits seamlessly. To evaluate our approach, we establish a benchmark and an evaluation protocol for part editing. Experiments show that our approach outperforms existing editing methods on all metrics and is preferred by users $66-90\%$ of the time in conducted user studies.
Part Editing, Image editing, Fine-grained editing, Diffusion Models
copyright: ccconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ; August 10–14, 2025; Vancouver, BC, Canadabooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers (SIGGRAPH Conference Papers ’25), August 10–14, 2025, Vancouver, BC, Canadadoi: 10.1145/3721238.3730747isbn: 979-8-4007-1540-2/2025/08submissionid: 1255ccs: Computing methodologies Computer visionccs: Computing methodologies Image manipulation
@@
Figure 1. Our approach, PartEdit, enables a wide range of fine-grained edits, allowing users to create highly customizable changes. The edits are seamless, precisely localized, and of high visual quality with no leakage into unedited regions.
1. Introduction
@@
Figure 2. A visualization for the cross-attention maps of SDXL (Podell et al., 2024) that corresponds to different words of the textual prompt. Object parts such as “head” and “hood” are not well-localized, indicating that the model lacks a sufficient understanding of these parts. DiT-based model analysis in Appendix R.
@@
Diffusion models (Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Podell et al., 2024; Esser et al., 2024) have significantly advanced image generation, achieving unprecedented levels of quality and fidelity. This progress is generally attributed to their large-scale training on the LAION-5B dataset (Schuhmann et al., 2022) with image-text pairs, leading to a profound understanding of images and their semantics. Recent image editing methods (Hertz et al., 2022; Brooks et al., 2023; Brack et al., 2024; Parmar et al., 2023; Tumanyan et al., 2023; Kawar et al., 2023; Huang et al., 2024; Andonian et al., 2023; Bar-Tal et al., 2022; Ju et al., 2024a; Chen et al., 2024a; Li et al., 2024; Lin et al., 2024; Chen et al., 2024b; Meng et al., 2021; Tang et al., 2024; Nichol et al., 2022) have capitalized on this understanding to perform a wide range of edits to enhance the creative capabilities of artists and designers. These methods allow users to specify desired edits through text prompts, enabling both semantic edits, such as modifying objects or their surroundings, and artistic adjustments, like changing style and texture. Ideally, these edits must align with the requested textual prompts while being seamlessly integrated and accurately localized in the image.
@@
In this paper, we address the limitations of pre-trained diffusion models in their understanding of object parts. By expanding their semantic knowledge, we enable fine-grained image edits, providing creators with greater control over their images. We achieve this by training part-specific tokens that specialize in localizing the editing region at each denoising step. Based on this localization, we develop feature blending and adaptive thresholding strategies that ensure seamless and high-quality editing while preserving the unedited areas. Our novel feature blending happens at each layer, at each timestep using nonbinary masks. To learn the part tokens, we design a token optimization process (Zhou et al., 2022) tailored for fine-grained editing, utilizing existing object part datasets (Donadello and Serafini, 2016; He et al., 2022) or user-provided datasets. This optimization process allows us to keep the pre-trained diffusion model frozen, thereby expanding its semantic understanding without compromising generation quality or existing knowledge. To evaluate our approach and to facilitate the development of future fine-grained editing approaches, we introduce a benchmark and an evaluation protocol for part editing. Experiments show that our approach outperforms all methods in comparison on all metrics and is preferred by users $66-90\%$ of the time in conducted user studies. Code and data for this paper are available at the project page https://gorluxor.github.io/part-edit/.
2. Text-to-Image Diffusion Models
Diffusion models are probabilistic generative models that attempt to learn an approximation of a data distribution $p(\mathbf{x})$. This is achieved by progressively adding noise to each data sample ${\mathbf{x}_{0}\sim p(\mathbf{x})}$ throughout $T$ timesteps until it converges to an isotropic Gaussian distribution as $T\rightarrow\infty$. During inference, a sampler such as DDIM (Song et al., 2020) is used to reverse this process starting from Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ that is iteratively denoised until we obtain a noise-free sample $\hat{\mathbf{x}}_{0}^{0}$. The reverse process for timestep $t\in[T,0]$ is computed as:
(1) $x_{t-1}=$ $\sqrt{\alpha_{t-1}}\ \hat{x}_{0}^{t}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\ \epsilon^{t}_{\theta}(x_{t})+\sigma_{t}\epsilon_{t}\enspace,$
@@ 2. Text-to-Image Diffusion Models
To generate an image that adheres to a user-provided input, e.g., a textual prompt $\mathcal{P}$, the model can be trained conditionally. In this setting, the UNet is conditioned on the prompt $\mathcal{P}$, and the noise prediction in Equation 1 is computed as $\epsilon^{t}_{\theta}(x_{t},\mathcal{P})$. More specifically, the textual prompt is embedded through a textual encoder, e.g., CLIP (Radford et al., 2021), to obtain an embedding $E\in\mathrm{R}^{77\times 2048}$ (for SDXL). This textual embedding interacts with the UNet image features $F$ within cross-attention modules at UNet block $i$ to compute attention as:
(2) $A_{i}=\text{{Softmax}}$ $\left(\dfrac{Q_{i}\ K_{i}^{\top}}{\sqrt{{d_{k_{i}}}}}\right)\enspace,$
@@
3. PartEdit: Fine-Grained Image Editing
Figure 3. An overview of our proposed approach for fine-grained part editing. For an object part $p$, we collect a dataset of images $\mathcal{I}^{p}$ and their corresponding part annotation masks $\mathcal{Y}^{p}$. To optimize a textual token to localize this part, we initialize a random textual embedding $\hat{E}$ that initially generates random cross-attention maps. During optimization, we invert images in $\mathcal{I}^{p}$ and optimize the part token so that the cross-attention maps at different layers and timesteps match the part masks in $\mathcal{Y}^{p}$. After optimizing the token, it can be used during inference to produce a localization mask at each denoising step. These localization masks are used to perform feature bending between the source and the edit image trajectories. Note that we visualize three instances of SDXL (Podell et al., 2024) for illustration, but in practice, this is done with the same model in a batch of three.
A common approach for editing images in diffusion-based methods involves manipulating the cross-attention maps (Hertz et al., 2022; Parmar et al., 2023; Epstein et al., 2023). If the cross-attention maps do not accurately capture the editing context, e.g., for editing object parts, these methods are likely to fail, as demonstrated in Figure 1. An intuitive solution to this problem is expanding the knowledge of the pre-trained diffusion models to understand object parts. This can be accomplished by fine-tuning the model with additional data of image/text pairs where the text is detailed. However, this approach can be costly due to the extensive annotation required for fine-tuning the model, and there is no guarantee that the model will automatically learn to identify object parts effectively from text.
An alternative approach is leveraging token optimization (Zhou et al., 2022) to learn new concepts through explicit supervision of cross-attention maps. This was proven successful in several applications (Hedlin et al., 2023; Khani et al., 2024; Marcos-Manchón et al., 2024) as it allows learning new concepts while keeping the model’s weights frozen. We leverage token optimization to perform fine-grained image editing of various object parts. We focus on optimizing tokens that can produce reliable non-binary blending masks at each diffusion step to localize the editing region. We supervise the optimization using either existing parts datasets such as PASCAL-Part (Donadello and Serafini, 2016), and PartImageNet (He et al., 2022) or a few user-annotated images.
3.1. Learning Part Tokens
@@
3.2. Choosing Timesteps and UNet Blocks to Optimize
Ideally, we would like to optimize over all timesteps and UNet layers. However, this is computationally and memory-expensive due to the large dimensionality of intermediate features in diffusion models, especially SDXL (Podell et al., 2024) that we employ. Therefore, we need to select a subset of layers and timesteps to achieve a good balance between localization accuracy and efficiency.
@@ 3.2. Choosing Timesteps and UNet Blocks to Optimize
Figure 4. Impact of timestep choice in the token optimization process. Intermediate timesteps achieve reasonable localization for both big and small parts.
3.3. Part Editing
@@
(5) $\mathcal{T}(X)$ =$=\begin{cases}1&\text{if }X\geq\omega\\ x&\text{if }k/2\leq X<\omega\\ 0,&X<k/2\\ \end{cases},$
@@
where OTSU is the OTSU thresholding (Otsu et al., 1975), $\omega$ is a tunable tolerance for the transition between the edited parts and the original object. We find that $\omega=3k/2$ achieves the best visual quality, and we fix it for all experiments. This criterion ensures suppressing the background noise and a smooth transition between the edited part and the rest of the object. Finally, we employ $M_{t}$ to blend the features between the source and the editing paths as:
(6) $\hat{F}_{i,t}^{e}=\mathcal{T}(M_{t})\ \hat{F}_{i,t}^{e}+(1-\mathcal{T}(M_{t}))\ \hat{F}_{i,t}^{s}$
@@
3.4. Real Image Editing
Our approach can also perform part edits on real images by incorporating a real image inversion method, e.g., Ledits++ (Brack et al., 2024) or EF-DDPM (Huberman-Spiegelglas et al., 2024). In this setting, the role of the inversion method is to estimate the diffusion trajectory $x_{0}\dots x_{T}$ for a given real image. This estimated trajectory is then used as the source path in Figure 3. To obtain the source prompt $\mathcal{P}^{s}$ of the real images, we use the image captioning approach, BLIP2 (Li et al., 2023), which is commonly used for this purpose. Finally, the edit is applied where the localization and editing paths are similar to the synthetic setting.
3.5. Discussion
@@
Figure 5. Qualitative comparison on synthetic images from the PartEdit benchmark. Our method outperforms both iP2P (Brooks et al., 2023) and P2P (Hertz et al., 2022) on the synthetic setting. Showcasing good localization while integrating seamlessly into the scene, illustrated by the third row with a formal torso edit.
@@
Figure 6. Qualitative comparison against EF-DDPM (Huberman-Spiegelglas et al., 2024) and Ledits++ (Brack et al., 2024) on real image editing.
@@
Figure 7. Visualization of editing 2 parts at the same time. Note that the attention maps showcase average cross-attention across all time steps.
4. Experiments
@@
In this section, we provide an evaluation of our proposed approach, a comparison against existing text-based editing approaches, and an ablation study. To facilitate these aspects, we create a manually annotated benchmark of few parts. For a comprehensive evaluation, we compare against synthetic and real image editing approaches.
4.1. Evaluation Setup
4.1.1. Synthetic Image Editing
@@
We base our method on a pre-trained SDXL (Podell et al., 2024) For sampling, we employ the DDIM sampler (Song et al., 2020) with $T=50$ denoising steps and default scheduling parameters as (Hertz et al., 2022). For token optimization, we train on 10-20 images per token for 2000 optimization steps. We set $t_{e}=T=50$ for complete preservation of the unedited regions. We compare against two popular text-based editing approaches: Prompt-to-Prompt (P2P) (Hertz et al., 2022) and Instruct-Pix2Pix (iP2P) (Brooks et al., 2023).
We use the SDXL implementation of P2P [^1] https://github.com/RoyiRa/prompt-to-prompt-with-sdxl, and the Diffusers (Wolf et al., 2020) implementation of iP2P [^2] https://huggingface.co/docs/diffusers/en/api/pipelines/pix2pix.
4.1.2. Real Image Editing
@@
We use Ledits++ (Brack et al., 2024) inversion method to invert the images where the source prompt $\mathcal{P}^{s}$ is produced by BLIP2 (Li et al., 2023) as described in Section 3.4. We compare against Ledits++ and EF-DDPM (Huberman-Spiegelglas et al., 2024) with the same base model of SD2.1 (Rombach et al., 2022).
4.1.3. PartEdit Benchmark
@@
We create a synthetic and real benchmark of 7 object parts: <humanoid-head>, <humanoid-torso>, <human -hair> <animal-head>, <car-body>, <car-hood>, and <chair-seat>. For the synthetic part, we generate random source prompts of the objects of interest at random locations and select random edits from pre-defined lists. We generate a total of 60 synthetic images and manually annotate the part of interest. For the real part, we collect 13 images from the internet and manually annotate and assign editing prompts to them. We denote the synthetic and real benchmarks as PartEdit-Synth and PartEdit-Real, respectively. More details are provided in the Appendix I.
4.1.4. Evaluation Metrics